Calculate the probability that the speed of a chlorine molecule will lie between 200 m s-1 and 220 m s-1

at 300 K. Given that mass of chlorine molecule = 70u, kB = 1.38 × 10-23JK-1 and NA = 6.02 × 1026 kmol

and NA= 6.02×1026 kmol⁻¹.


1
Expert's answer
2021-04-16T07:25:21-0400

"\frac{\Delta n}{n}=4\pi(\frac{m}{2\pi k T})^{\frac 32}v^2e^{-\frac{mv^2}{2kT}}\Delta v=0.054."